# Profinite rigidity witnessed by Dehn fillings of cusped hyperbolic 3-manifolds

### 03 Mar 2026 v3

@@ 1.1. Profinite rigidity from Dehn filling
    The Berge manifold, i.e. the complement of the link $L_{A}$ in $S^{3}$ (Figure 5(a)).
@@ 1.2. Application to knot complements
In fact, Theorem A, i.e. the profinite detection of Dehn filling, shows that the profinite completion of fundamental group seems to impose strong restrictions on a cusped hyperbolic 3-manifold. This may provide more confidence in the following conjecture.
@@ 2.5. Peripheral structure
According to [28] any abelian subgroup of $\pi_{1}M$ is separable. As a result, $\pi_{1}M$ induces the full profinite topology on $\pi_{1}\partial_{i}M\cong\mathbb{Z}^{2}$. The standard theories of profinite groups [41, Lemma 3.2.6] implies that $\widehat{\pi_{1}\partial_{i}M}\to\widehat{\pi_{1}M}$ is injective. Note that the image of this homomorphism is exactly $\overline{\pi_{1}\partial_{i}M}$. Thus, $\overline{\pi_{1}\partial_{i}M}\cong\widehat{\pi_{1}\partial_{i}M}\cong\widehat{\mathbb{Z}}^{2}$. ∎
@@ 3.2. Dehn filling
Now we prove that $f_{\star}$ respects the peripheral structure. Indeed, $\partial M_{(c_{i})}$ and $\partial N_{(\overline{\psi}(c_{i}))}$ consists of the boundary components $\partial_{j}M$ and $\partial_{j}N$, where $c_{j}=\varnothing$. The corresponding peripheral subgroups are the images of $\iota_{j_{\star}}:\pi_{1}\partial_{j}M\subseteq\pi_{1}M\to\pi_{1}M/\langle\!\langle\gamma_{1},\cdots,\gamma_{n}\rangle\!\rangle\cong\pi_{1}M_{(c_{i})}$ and $\rho_{j_{\ast}}:\pi_{1}\partial_{j}N\subseteq\pi_{1}N\to\pi_{1}N/\langle\!\langle\psi_{1}(\gamma_{1}),\cdots,\psi_{n}(\gamma_{n})\rangle\!\rangle\cong\pi_{1}N_{(\overline{\psi}(c_{i}))}$ respectively. Recall that $f$ respects the peripheral structure, i.e. $f(\overline{\pi_{1}\partial_{j}M})=g_{j}\cdot\overline{\pi_{1}\partial_{j}N}\cdot g_{j}^{-1}$ for some $g_{j}\in\widehat{\pi_{1}N}$. Desending to $f_{\star}:\widehat{\pi_{1}M_{(c_{i})}}\cong\widehat{\pi_{1}N_{(\overline{\psi}_{i}(c_{i}))}}$, we have $f_{\star}(\overline{\iota_{j_{\star}}(\pi_{1}\partial_{j}M)})=\phi(g_{j})\cdot\overline{\rho_{j_{\star}}(\pi_{1}\partial_{j}N)}\cdot\phi(g_{j})^{-1}$, where $\phi:\widehat{\pi_{1}N}\to\widehat{\pi_{1}N}/\overline{\langle\!\langle\psi_{1}(\gamma_{1}),\cdots,\psi_{n}(\gamma_{n})\rangle\!\rangle}\cong\widehat{\pi_{1}N_{(\overline{\psi}(c_{i}))}}$ denotes the quotient homomorphism. Thus, $f_{\star}$ also respects the peripheral structure. ∎
@@ 4.1. Detection of knot and link complements
Let $M=S^{3}\setminus\mathring{N}(L)$ be the complement of a hyperbolic link $L$ in $S^{3}$ (i.e. the link complement is hyperbolic). Suppose $N$ is a compact orientable 3-manifold such that $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$, then $N$ is homeomorphic to the complement of a hyperbolic link $L^{\prime}\subseteq S^{3}$, where $L$ and $L^{\prime}$ have the same number of components.
@@
In particular, when $L$ has exactly one component, i.e. $L$ is a knot, we derive that any compact orientable 3-manifold profinitely isomorphic to a hyperbolic knot complement is also a hyperbolic knot complement.
@@ 4.2. Identifying Dehn surgery coefficients
Let $(m,l)$ and $(m^{\prime},l^{\prime})$ be the oriented meridian-longitude pair of $K$ and $K^{\prime}$ respectively. Note that the longitude is the rational null-homologous boundary slope of a knot complement. Thus, according to Lemma 4.4 (2), $\overline{\psi}([l])=[l^{\prime}]$, i.e. $\psi(l)=\pm l^{\prime}$.
Now we show that $\psi(m)=\pm m^{\prime}$. Indeed, $M_{[m]}\cong S^{3}$, and $\widehat{\pi_{1}M^{\prime}_{\overline{\psi}[m]}}\cong\widehat{\pi_{1}M_{[m]}}$ is the trivial group. The same argument as in the proof of Theorem 4.1, i.e. the residual finiteness of the fundamental group together with the validity of Poincaré conjecture, implies that $M^{\prime}_{\overline{\psi}[m]}$ is homeomorphic to $S^{3}$. Note that $K^{\prime}$ is a non-trivial knot; according to [24, Theorem 2], the only Dehn-surgery of $K^{\prime}$ yielding $S^{3}$ is the trivial one, i.e. the $\infty$-surgery. Therefore, $\overline{\psi}[m]$ represents the same boundary slope as $m^{\prime}]$, i.e. $\psi(m)=\pm m^{\prime}$.
@@ 4.3. Criteria for profinite rigidity of a knot complement
Suppose $J$ is a knot in $S^{3}$, and there is a homeomorphism $F:J(0)\to K(0)$.
Case 1: $F$ is orientation-preserving.
In this case, by Definition 4.7 (2), $J$ is isotopic to $K$.
Case 2: $F$ is orientation-reversing.
Let $\overline{J}$ denote the mirror image of $J$. Then $\overline{J}(0)\cong J(0)$ by an orientation-reversing homeomorphism. Thus, there exists an orientation-preserving homeomorphism $\overline{J}(0)\cong K(0)$. Then, Definition 4.7 (2) implies that $\overline{J}$ is isotopic to $K$, i.e. $J$ is the mirror image of $K$. ∎
@@ 5. Profinite detection of exceptional Dehn fillings
as the collection of all exceptional slopes on $\partial_{i_{0}}M$. Indeed, [45, Theorem 5.8.2] implies that only finitely many slopes on $\partial_{i_{0}}M$ yield exceptional Dehn fillings, i.e. $\mathcal{E}(M,\partial_{i_{0}}M)<+\infty$.
@@ 5.2. Detecting toroidality
Note that $M$ is toroidal if and only if one of $M_{1},\cdots,M_{s}$ is toroidal, and $N$ is toroidal if and only if one of $N_{1},\cdots,N_{s^{\prime}}$ is toroidal. Thus, it suffices to consider each irreducible and boundary-incompressible prime factor, and show that $M_{i}$ is toroidal if and only if $N_{i}$ is toroidal.
Based on the hyperbolization theorem [3, Theorem 1.7.5], $M_{i}$ is toroidal if and only if the following three conditions hold simultaneously.
@@
    $M_{i}$ is not a finite-volume hyperbolic manifold;
@@
    $M_{i}$ is not a closed small Seifert fibered space, i.e. $M_{i}$ does not admit a Seifert fibration over a sphere with at most 3 singular points;
@@
    $M_{i}$ is not homeomorphic to any of the following listed Seifert fibered spaces with boundary:
@@
    • $P\times S^{1}=(P;)$, where $P$ denotes a pair of pants;
    • $(S^{1}\times I;(p,q))$, $p>1$, i.e. a Seifert fibered manifold with base orbifold an annulus with one singular point;
    • $(D^{2};(p_{1},q_{1}),(p_{2},q_{2}))$, $p_{1},p_{2}>1$ and $(p_{1},p_{2})\neq(2,2)$, i.e. a Seifert fibered manifold with base orbifold a disk with two singular points and having negative Euler characteristic.
@@
Moving on to condition (ii), it suffices to show that when $M_{i}$ is a closed small Seifert fibered space, $N_{i}$ is also a closed small Seifert fibered space. Indeed, $N_{i}$ is also closed since $f_{i}$ respects the peripheral structure, and it follows from [53, Theorem 8.4] that $N_{i}$ is also a Seifert fibered space. When $\pi_{1}M_{i}$ is finite, the residual finiteness of $\pi_{1}N_{i}$ (cf. Proposition 2.2) implies that $\pi_{1}N_{i}$ is also finite, and $N_{i}$ is hence a small Seifert fibered space admitting $\mathbb{S}^{3}$ geometry. When $\pi_{1}M_{i}$ is infinite, [47, Theorem 1.2] implies that either $M_{i}$ and $N_{i}$ are homeomorphic, or $M_{i}$ and $N_{i}$ both admit $\mathbb{H}^{2}\times\mathbb{R}$ geometry and share the same base orbifold. In either case, $N_{i}$ is also a closed small Seifert fibered space.
For (iii), it suffices to show that when $M_{i}$ belongs to this list, $N_{i}$ also belongs to the list. Since $M_{i}$ and $N_{i}$ are irreducible with non-empty incompressible toral boundary, it follows from a combination of [55] and [54] (see also [56, Theorem 5.5]) that $N_{i}$ is also Seifert fibered. Therefore, $f_{i}:\widehat{\pi_{1}M_{i}}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N_{i}}$ is a profinite isomorphism between Seifert fibered spaces with boundary, which respects the peripheral structure. [47, Theorem 1.3] then implies that $M_{i}$ and $N_{i}$ have the same base orbifolds. The list of Seifert fibered spaces in (iii) are only characterized by their base orbifolds. Therefore, when $M_{i}$ belongs to this list, so does $N_{i}$, which finishes the proof. ∎
@@ 5.3. Detection of embedded Klein bottles
    $M$ is closed, admitting $Sol$-geometry, but does not fiber over $S^{1}$.
@@
Now let us consider the case when $M$ has non-empty boundary. Then the above reasoning shows that $N$ also has non-empty boundary. If $M$ is homeomorphic to $S^{1}\times S^{1}\times I$ or $\mathcal{N}$, then it is easy to verify that $N\cong M$ (cf. [56, Proposition 8.2]), and the conclusion follows directly. Thus, we may assume that $M$ fibers over a hyperbolic orbifold. In this case, $M$ admits a unique Seifert fibration, and we denote the base orbifold as $\mathcal{O}_{M}$. In addition, $N$ also fibers over a hyperbolic orbifold, and we denote the unique base orbifold as $\mathcal{O}_{N}$. In this case, the base orbifolds $\mathcal{O}_{M}$ and $\mathcal{O}_{N}$ may not be identical. However, in [56, Proposition 8.2], it is proven that the isomorphism type of $\widehat{\pi_{1}M}$ (resp. $\widehat{\pi_{1}N}$) determines whether the base orbifold $\mathcal{O}_{M}$ (resp. $\mathcal{O}_{N}$) is orientable, and determines the isomorphism type of the orbifold fundamental group $\pi_{1}(\mathcal{O}_{M})$ (resp. $\pi_{1}(\mathcal{O}_{N})$). Thus, $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$ implies that $\mathcal{O}_{M}$ is non-orientable if and only if $\mathcal{O}_{N}$ is non-orientable. In addition, there is a one-to-one correspondence between the conjugacy classes of maximal torsion subgroups in the orbifold fundamental group with the cone points of a compact 2-orbifold with discrete singular set and non-empty boundary. Therefore, $\mathcal{O}_{M}$ contains two singular points of index $2$ if and only if $\pi_{1}(\mathcal{O}_{M})\cong\pi_{1}(\mathcal{O}_{N})$ contains two conjugacy classes of maximal torsion subgroups isomorphic to $\mathbb{Z}/2\mathbb{Z}$, which implies that $\mathcal{O}_{N}$ also contains two singular points of index $2$. ∎
@@
First, we suppose that $M$ satisfies condition (i) in Lemma 5.7, i.e. $M$ is a closed $Sol$-manifold which is not fibered over $S^{1}$. Since $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$, it follows from [40, Corollary 4.12] that $N$ is also closed, and [53, Theorem 8.4] shows that $N$ also admits $Sol$-geometry. In addition, [31, Corollary 1.2] shows that the profinite completion of fundamental group of a compact, orientable 3-manifold determines whether this 3-manifold fibers over $S^{1}$. Thus, $N$ is also non-fibered, and satisfies condition (i) of Lemma 5.7.
Next, we suppose that $M$ satisfies condition (ii) in Lemma 5.7. Similarly, [53, Theorem 8.4] implies that $N$ is not a closed $Sol$-manifold. [55, Theorem B] shows that the profinite completion of fundamental group detects the JSJ-decomposition of a compact, orientable, irreducible 3-manifold with empty or incompressible toral boundary which does not admit $Sol$-geometry. In fact, [55, Theorem B] was originally proven for closed manifolds, but the proof can be easily generalized to the bounded case, see for example [56, Theorem 5.5]. In particular, there is a one-to-one correspondence between the Seifert fibered JSJ-pieces of $M$ and $N$, such that the corresponding pieces are profinitely isomorphic. Since one of the Seifert pieces of $M$ has a base orbifold which is either non-orientable or contains two singular points of index $2$, and this Seifert piece is obviously boundary incompressible and aspherical, it follows from Lemma 5.8 that the corresponding Seifert piece of $N$ also has a base orbifold which is either non-orientable or contains two singular points of index $2$. Therefore, $N$ also satisfies condition (ii) in Lemma 5.7. ∎

### 16 Mar 2026 v4

@@ 1.1. Profinite rigidity from Dehn filling
    The Berge manifold, ie the complement of the link $L_{A}$ in $S^{3}$ (Figure 5(a)).
@@ 1.2. Application to knot complements
In fact, Theorem A, ie the profinite detection of Dehn filling, shows that the profinite completion of fundamental group seems to impose strong restrictions on a cusped hyperbolic 3-manifold. This may provide more confidence in the following conjecture.
@@ 2.5. Peripheral structure
According to [28], any abelian subgroup of $\pi_{1}M$ is separable. As a result, $\pi_{1}M$ induces the full profinite topology on $\pi_{1}\partial_{i}M\cong\mathbb{Z}^{2}$. The standard theories of profinite groups [41, Lemma 3.2.6] implies that $\widehat{\pi_{1}\partial_{i}M}\to\widehat{\pi_{1}M}$ is injective. Note that the image of this homomorphism is exactly $\overline{\pi_{1}\partial_{i}M}$. Thus, $\overline{\pi_{1}\partial_{i}M}\cong\widehat{\pi_{1}\partial_{i}M}\cong\widehat{\mathbb{Z}}^{2}$. ∎
@@ 3.2. Dehn filling
Now we prove that $f_{\star}$ respects the peripheral structure. Indeed, $\partial M_{(c_{i})}$ and $\partial N_{(\overline{\psi}(c_{i}))}$ consist of the boundary components $\partial_{j}M$ and $\partial_{j}N$, where $c_{j}=\varnothing$. The corresponding peripheral subgroups are the images of $\iota_{j_{\star}}:\pi_{1}\partial_{j}M\hookrightarrow\pi_{1}M\to\pi_{1}M/\langle\!\langle\gamma_{1},\cdots,\gamma_{n}\rangle\!\rangle\cong\pi_{1}M_{(c_{i})}$ and $\rho_{j_{\ast}}:\pi_{1}\partial_{j}N\hookrightarrow\pi_{1}N\to\pi_{1}N/\langle\!\langle\psi_{1}(\gamma_{1}),\cdots,\psi_{n}(\gamma_{n})\rangle\!\rangle\cong\pi_{1}N_{(\overline{\psi}(c_{i}))}$ respectively. Recall that $f$ respects the peripheral structure, ie $f(\overline{\pi_{1}\partial_{j}M})=g_{j}\cdot\overline{\pi_{1}\partial_{j}N}\cdot g_{j}^{-1}$ for some $g_{j}\in\widehat{\pi_{1}N}$. Desending to $f_{\star}:\widehat{\pi_{1}M_{(c_{i})}}\cong\widehat{\pi_{1}N_{(\overline{\psi}_{i}(c_{i}))}}$, we have $f_{\star}(\overline{\iota_{j_{\star}}(\pi_{1}\partial_{j}M)})=\phi(g_{j})\cdot\overline{\rho_{j_{\star}}(\pi_{1}\partial_{j}N)}\cdot\phi(g_{j})^{-1}$, where $\phi:\widehat{\pi_{1}N}\to\widehat{\pi_{1}N}/\overline{\langle\!\langle\psi_{1}(\gamma_{1}),\cdots,\psi_{n}(\gamma_{n})\rangle\!\rangle}\cong\widehat{\pi_{1}N_{(\overline{\psi}(c_{i}))}}$ denotes the quotient homomorphism. Thus, $f_{\star}$ also respects the peripheral structure. ∎
@@ 4.1. Detection of knot and link complements
Let $M=S^{3}\setminus\mathring{N}(L)$ be the complement of a hyperbolic link $L$ in $S^{3}$ (ie the link complement is hyperbolic). Suppose $N$ is a compact orientable 3-manifold such that $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$, then $N$ is homeomorphic to the complement of a hyperbolic link $L^{\prime}\subseteq S^{3}$, where $L$ and $L^{\prime}$ have the same number of components.
@@
In particular, when $L$ has exactly one component, ie $L$ is a knot, we derive that any compact orientable 3-manifold profinitely isomorphic to a hyperbolic knot complement is also a hyperbolic knot complement.
@@ 4.2. Identifying Dehn surgery coefficients
Let $(m,l)$ and $(m^{\prime},l^{\prime})$ be the oriented meridian-longitude pair of $K$ and $K^{\prime}$ respectively. Note that the longitude is the rational null-homologous boundary slope of a knot complement. Thus, according to Lemma 4.4 (2), $\overline{\psi}([l])=[l^{\prime}]$, ie $\psi(l)=\pm l^{\prime}$.
Now we show that $\psi(m)=\pm m^{\prime}$. Indeed, $M_{[m]}\cong S^{3}$, and $\widehat{\pi_{1}M^{\prime}_{\overline{\psi}[m]}}\cong\widehat{\pi_{1}M_{[m]}}$ is the trivial group. The same argument as in the proof of Theorem 4.1, ie the residual finiteness of the fundamental group together with the validity of Poincaré conjecture, implies that $M^{\prime}_{\overline{\psi}[m]}$ is homeomorphic to $S^{3}$. Note that $K^{\prime}$ is a non-trivial knot; according to [24, Theorem 2], the only Dehn-surgery of $K^{\prime}$ yielding $S^{3}$ is the trivial one, ie the $\infty$-surgery. Therefore, $\overline{\psi}[m]$ represents the same boundary slope as $m^{\prime}]$, ie $\psi(m)=\pm m^{\prime}$.
@@ 4.3. Criteria for profinite rigidity of a knot complement
Suppose $J$ is a knot in $S^{3}$, and there is a homeomorphism $F:J(0)\to K(0)$. There are two possibilities. First, if $F$ is orientation-preserving, then Definition 4.7 (2) implies that $J$ is isotopic to $K$. Second, if $F$ is orientation-reversing, let $\overline{J}$ denote the mirror image of $J$. Then $\overline{J}(0)\cong J(0)$ by an orientation-reversing homeomorphism. Thus, there exists an orientation-preserving homeomorphism $\overline{J}(0)\cong K(0)$. Then, Definition 4.7 (2) implies that $\overline{J}$ is isotopic to $K$, ie $J$ is the mirror image of $K$. ∎
@@ 5. Profinite detection of exceptional Dehn fillings
as the collection of all exceptional slopes on $\partial_{i_{0}}M$. Indeed, [45, Theorem 5.8.2] implies that only finitely many slopes on $\partial_{i_{0}}M$ yield exceptional Dehn fillings, ie $\mathcal{E}(M,\partial_{i_{0}}M)<+\infty$.
@@ 5.2. Detecting toroidality
Note that $M$ is toroidal if and only if one of $M_{1},\cdots,M_{s}$ is toroidal, and $N$ is toroidal if and only if one of $N_{1},\cdots,N_{s^{\prime}}$ is toroidal. Thus, it suffices to show that for each irreducible and $\partial$-irreducible prime factor, $M_{i}$ is toroidal if and only if $N_{i}$ is toroidal.
Based on the hyperbolization theorem (see [3, Theorem 1.7.5]) and standard results on Seifert fibered spaces, $M_{i}$ is not toroidal if and only if $M_{i}$ falls into one of the following 3 possibilities.
@@
    $M_{i}$ is a finite-volume hyperbolic manifold;
@@
    $M_{i}$ is a closed Seifert fibered space that admits a Seifert fibration over $S^{2}$ with at most 3 exceptional fibers, but does not admit $\mathbb{E}^{3}$ geometry;
@@
    $M_{i}$ is homeomorphic to one of the following listed Seifert fibered spaces with boundary:
@@
    • $P\times S^{1}=(P;)$, where $P=\Sigma_{0,3}$ denotes a pair of pants;
    • $(S^{1}\times I;(p,q))$, $p>1$, ie a Seifert fibered manifold with base orbifold an annulus with one singular point;
    • $(D^{2};(p_{1},q_{1}),(p_{2},q_{2}))$, $p_{1},p_{2}>1$ and $(p_{1},p_{2})\neq(2,2)$, ie a Seifert fibered manifold with base orbifold a disk with two singular points and having negative Euler characteristic.
@@
Moving on to condition (ii), suppose $M_{i}$ is a closed Seifert fibered space that admits a Seifert fibration over $S^{2}$ with at most 3 exceptional fibers, and $M_{i}$ does not admit $\mathbb{E}^{3}$ geometry. Then $N_{i}$ is also closed since $f_{i}$ respects the peripheral structure, and it follows from [53, Theorem 8.4] that $N_{i}$ is also a Seifert fibered space admitting the same geometry of $M_{i}$. In particular, $N_{i}$ does not admit $\mathbb{E}^{3}$ geometry. In addition, $M_{i}$ and $N_{i}$ cannot admit $\mathbb{S}^{2}\times\mathbb{E}^{1}$ geometry since they are irreducible. If $M_{i}$ admits $\mathbb{S}^{3}$ geometry, then $N_{i}$ also admits $\mathbb{S}^{3}$ geometry, and any such manifold admits a Seifert fibration over $S^{2}$ with at most 3 exceptional fibers. If they admit $Nil$ geometry or $\widetilde{\mathrm{SL}(2,\mathbb{R})}$ geometry, [47, Theorem 1.2] implies that $M_{i}$ is homeomorphic with $N_{i}$, so the condition holds automatically for $N_{i}$. Finally, if $M_{i}$ and $N_{i}$ admit $\mathbb{H}^{2}\times\mathbb{E}^{1}$ geometry, then they admit unique Seifert fibrations; and by [47, Theorem 5.2], $M_{i}$ and $N_{i}$ have the same base orbifold though they are not necessarily homeomorphic. Hence, $N_{i}$ also admits a Seifert fibration over $S^{2}$ with at most 3 exceptional fibers if $M_{i}$ does.
For (iii), suppose $M_{i}$ belongs to this list. Since $M_{i}$ and $N_{i}$ are irreducible with non-empty incompressible toral boundary, it follows from the results of Wilton-Zalesskii [55, 54] that $N_{i}$ is also Seifert fibered. The isomorphism $f_{i}:\widehat{\pi_{1}M_{i}}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N_{i}}$ respects the peripheral structure, so [47, Theorem 5.8] implies that $M_{i}$ and $N_{i}$ have the same base orbifolds. The list of Seifert fibered spaces in (iii) are only characterized by their base orbifolds. Therefore, $N_{i}$ belongs to the list when $M_{i}$ does, which finishes the proof. ∎
@@ 5.3. Detection of embedded Klein bottles
    $M$ is closed, admitting $Sol$ geometry, but does not fiber over $S^{1}$.
@@
Now let us consider the case when $M$ has non-empty boundary. Then the above reasoning shows that $N$ also has non-empty boundary. If $M$ is homeomorphic to $S^{1}\times S^{1}\times I$ or $\mathcal{N}$, then it is easy to verify that $N\cong M$ (cf. [56, Proposition 8.2]), and the conclusion follows directly. Thus, we may assume that $M$ fibers over a hyperbolic orbifold. In this case, $M$ admits a unique Seifert fibration, and we denote the base orbifold as $\mathcal{O}_{M}$. In addition, $N$ also fibers over a hyperbolic orbifold, and we denote the unique base orbifold as $\mathcal{O}_{N}$. In this case, the base orbifolds $\mathcal{O}_{M}$ and $\mathcal{O}_{N}$ may not be identical. However, in [56, Proposition 8.2], it is proven that the isomorphism type of $\widehat{\pi_{1}M}$ (resp. $\widehat{\pi_{1}N}$) determines whether the base orbifold $\mathcal{O}_{M}$ (resp. $\mathcal{O}_{N}$) is orientable, and determines the isomorphism type of the orbifold fundamental group $\pi_{1}(\mathcal{O}_{M})$ (resp. $\pi_{1}(\mathcal{O}_{N})$). Thus, $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$ implies that $\mathcal{O}_{M}$ is non-orientable if and only if $\mathcal{O}_{N}$ is non-orientable. In addition, there is a one-to-one correspondence between the conjugacy classes of maximal torsion subgroups in $\pi_{1}(\mathcal{O}_{M})$ (resp. $\pi_{1}(\mathcal{O}_{N})$) with the cone points in $\mathcal{O}_{M}$ and $\mathcal{O}_{N}$. To be explicit, $\mathcal{O}_{M}$ contains two singular points of index $2$ if and only if $\pi_{1}(\mathcal{O}_{M})\cong\pi_{1}(\mathcal{O}_{N})$ contains two conjugacy classes of maximal torsion subgroups isomorphic to $\mathbb{Z}/2\mathbb{Z}$, which implies that $\mathcal{O}_{N}$ also contains two singular points of index $2$. ∎
@@
First, we suppose that $M$ satisfies condition (i) in Lemma 5.7, ie $M$ is a closed $Sol$-manifold which is not fibered over $S^{1}$. Since $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$, it follows from [40, Corollary 4.12] that $N$ is also closed, and [53, Theorem 8.4] shows that $N$ also admits $Sol$-geometry. In addition, [31, Corollary 1.2] shows that the profinite completion of fundamental group of a compact, orientable 3-manifold determines whether this 3-manifold fibers over $S^{1}$. Thus, $N$ is also non-fibered, and satisfies condition (i) of Lemma 5.7.
Next, we suppose that $M$ satisfies condition (ii) in Lemma 5.7. Similarly, [53, Theorem 8.4] implies that $N$ is not a closed $Sol$-manifold. [55, Theorem B] shows that the profinite completion of the fundamental group detects the JSJ-decomposition of a compact, orientable, irreducible 3-manifold with empty or incompressible toral boundary which does not admit $Sol$ geometry. In fact, [55, Theorem B] was originally proven for closed manifolds, but the proof can be easily generalized to the bounded case, see for example [56, Theorem 5.5]. In particular, there is a one-to-one correspondence between the Seifert fibered JSJ-pieces of $M$ and $N$, such that the corresponding pieces are profinitely isomorphic. Since one of the Seifert pieces of $M$ has a base orbifold which is either non-orientable or contains two singular points of index $2$, and this Seifert piece is obviously boundary incompressible and aspherical, it follows from Lemma 5.8 that the corresponding Seifert piece of $N$ also has a base orbifold which is either non-orientable or contains two singular points of index $2$. Therefore, $N$ also satisfies condition (ii) in Lemma 5.7. ∎